# Bridged Semantic Alignment for Zero-shot 3D Medical Image Diagnosis11footnotemark: 1

### v1

@@ 4.1 Materials
We conduct experiments on two publicly available datasets: CT-RATE and RAD-ChestCT. The CT-RATE dataset consists of 3D CT volumes paired with medical reports, providing a valuable resource for evaluating cross-modal alignment methods in the medical domain. The RAD-ChestCT dataset includes 3D CT volumes with detailed annotations of various abnormalities, enabling the assessment of zero-shot diagnosis performance. Detailed descriptions of the datasets are as follows:
CT-RATE The CT-RATE dataset comprises 25,692 non-contrast chest CT volumes, expanded to 50,188 volumes through various reconstruction techniques, originating from 21,304 unique patients. Each CT volume is paired with a corresponding radiology report. A total of 47,149 CT volumes with corresponding reports are used for training, while 3,039 CT volumes are reserved for internal validation. The dataset includes 18 abnormality labels, extracted from reports using an automated algorithm. To further evaluate the zero-shot detection capabilities of our method, we curate 15 underrepresented abnormality labels from the validation set by GPT-4 [1] [^2] https://openai.com/index/gpt-4/, reflecting a long-tail distribution, to construct an extended CT-RATE-LT dataset. All curated labels are manually corrected by a radiologist to ensure annotation accuracy. The frequency of abnormalities in CT-RATE-LT is shown in Fig. 4.
RAD-ChestCT The RAD-ChestCT dataset contains 3,630 CT volumes, uniformly reconstructed using a single technique. This dataset includes 84 abnormality labels. As part of our external validation, as shown in Fig. 5, 27 abnormalities from RAD-ChestCT are mapped to the 18 abnormalities in CT-RATE to ensure consistency. To further demonstrate the strength of our method, we refer to the remaining 56 abnormalities from RAD-ChestCT as RAD-ChestCT-LT and evaluate the zero-shot diagnosis performance on these additional abnormalities. Note that the abnormality, “other_path”, is excluded from our analysis due to its lack of meaningful semantic descriptions.
@@ 4.2 Dataset Preprocessing
For the image data, we unify the spacing of CT volumes to (1.5 $\mathrm{mm}$, 1.5 $\mathrm{mm}$, 3 $\mathrm{mm}$), to ensure uniform resolution across the datasets. Random cropping is employed during the training stage to generate input volumes with a shape of $(224,224,112)$, which balances computational efficiency and sufficient spatial coverage for 3D images. The patch size in the ViT-base model is set to $(16,16,8)$, following the adaptation for 3D inputs. To enhance feature representation, CT images are clipped to a Hounsfield Unit (HU) range of $[-1000,1000]$, which focuses on clinically relevant tissues, and normalized to the pixel intensity range $[-1,1]$.
For the text data, we process both the original reports and the semantic summaries generated by GPT-4, splitting them into shorter sentences for analysis. As part of our data augmentation strategy, five sentences are randomly sampled and concatenated into a longer sentence during training, enabling diverse text representations that better capture the variability in medical reporting styles. This augmented sentence serves as the corresponding image report, improving the robustness of vision-language alignment.
Table 1: Zero-shot abnormality diagnosis performance comparison across internal (CT-RATE) and external (RAD-ChestCT) validation benchmarks, including shape dimensions $[H,W,D]$ and voxel spacing (mm). The best performance is highlighted as bold.
Method Setting Shape Spacing CT-RATE RAD-ChestCT
$[H,W,D]$ (mm) AUC ACC F1 score Precision AUC ACC F1 score Precision
CT-Net Supervised $[420,420,402]$- 60.3 58.1 63.1 23.9 54.4 54.0 58.7 28.5
VocabFine Fine-tuning $[480,480,224]$ $[0.75,0.75,1.5]$ 75.6 70.5 73.8 35.3 65.0 61.5 65.9 34.9
ClassFine Fine-tuning $[480,480,224]$ $[0.75,0.75,1.5]$ 75.6 68.9 72.4 33.9 64.3 60.7 64.9 34.9
CT-CLIP Zero-shot $[480,480,224]$ $[0.75,0.75,1.5]$ 73.1 66.8 70.7 32.3 62.9 59.5 64.2 33.6
BIUD Zero-shot $[320,224,64]$- 71.3 68.1 71.6 33.8 62.9 60.6 65.2 33.7
Merlin Zero-shot $[224,224,160]$ $[1.50,1.50,3.0]$ 72.8 67.2 70.9 33.7 64.4 61.9 66.3 34.8
fLVM Zero-shot $[480,352,224]$ $[0.75,0.75,1.5]$ 77.8 71.8 75.1 37.9 68.0 64.7 68.8 37.4
BrgSA Zero-shot $[224,224,112]$ $[1.50,1.50,3.0]$ 79.2 73.3 76.2 38.5 70.0 65.5 69.3 39.1
@@ 4.3 Implementation Details
In our experiments, we adapt ViT-B/16 as the image encoder by replacing the 2D convolutional layers in patch embedding module with 3D convolutional layers, resulting in a 3D ViT-B/16 architecture. For the text encoder, we fine-tune PubMedBERT on the text data of the CT-RATE dataset to better capture the linguistic characteristics and representation requirements of the medical domain. Furthermore, we utilize M3AE [7] for pretraining the image and text encoders. M3AE employs a masked cross-modal reconstruction strategy, where masked regions in one modality are reconstructed using information from the other modality. This approach enhances the semantic consistency between image and text features and strengthens their alignment for downstream tasks.
In our experiments, $\alpha$, $\beta$, and $\gamma$ are empirically set to 0.5, 1, and 1, respectively, based on preliminary trials to balance the contributions of different loss components. $K$ is set to 2048, as determined through the experimental search results presented later. We use the Adam optimizer with a learning rate of $5\times 10^{-5}$. The batch size is set to 8 for all experiments. The code is implemented using the PyTorch framework, and all experiments are conducted on a single NVIDIA A800 GPU. During the inference phase, we set the prompt template as “There is [abnormality]” for zero-shot abnormality diagnosis.
@@ 4.4 Evaluation Metrics
For the multi-label classification, we adapt Area under the ROC Curve (AUC), Accuracy (ACC), F1 score, and Precision as metrics for evaluating zero-shot abnormality diagnosis. For volume-to-volume retrieval task, we use Mean Average Precision at $Q$ (MAP@$Q$) metrics, where $Q$ represents the number of top-ranked retrieved volumes considered. In our experiment, $Q$ is set to $\{5,10,50\}$. For the report-to-volume retrieval task, we use Recall@$P$, where $P$ represents the number of retrieved images, to measure the proportion of correctly matched images within the top-$P$ retrieved results for a given query text. In our experiment, $P$ is set to $\{5,10,50,100\}$.
Table 2: Performance comparison for volume-to-volume retrieval task on CT-RATE-LT and RAD-ChestCT-LT datasets.
Dataset Method AUC ACC F1 score Precision
CT-RATE-LT VocabFine 69.2 71.2 79.5 2.2
CT-CLIP 76.9 78.1 86.6 1.0
Merlin 75.3 80.8 86.4 2.8
BrgSA 85.6 88.5 93.3 4.3
RAD-ChestCT-LT VocabFine 59.0 57.5 66.6 13.4
CT-CLIP 58.2 56.9 66.2 13.1
Merlin 59.1 57.0 66.1 13.6
BrgSA 62.9 59.9 68.4 14.6
@@ 4.5 Comparison with State-of-the-art Methods
We compare our method against four zero-shot approaches, supervised learning, and fine-tuning methods, as summarized in Table 1. The evaluation is conducted on both internal (CT-RATE) and external (RAD-ChestCT) validation benchmarks for zero-shot abnormality diagnosis. In addition, we examine the input image shape and spacing used by each method. Note that when images are resized to a fixed shape without maintaining consistent resolution, the spacing becomes undefined. To provide a comprehensive evaluation, we detail the comparison methods as follows:
Supervised Learning Method:
  • 1.
    CT-Net [10]: This method uses a deep convolutional neural network with transfer learning and 3D convolutions to perform multi-disease classification.
Fine-tuning Methods:
  • 1.
    VocabFine [15]: This method fine-tunes only the projection layers of the image and text encoders in CT-CLIP while freezing all other pre-trained layers, preserving open-vocabulary capabilities.
  • 2.
    ClassFine [15]: This method adds and trains a new classification layer on CT-CLIP while optionally freezing the pre-trained layers to retain existing feature representations.
Zero-shot Methods:
  • 1.
    CT-CLIP [15]: This method employs a contrastive vision-language pretraining approach to align 3D chest CT volumes with radiology reports, utilizing a vision transformer for 3D CT and a text transformer for embedding generation.
  • 2.
    BIUD [6]: This method aligns 3D CT images with 2D X-ray images through a language-guided retrieval strategy and enhances alignment using robust contrastive learning and entity-focused masking.
  • 3.
    Merlin [4]: This method employs multi-task learning to process 3D abdominal CT volumes, by integrating structured electronic health record data and unstructured radiology reports, optimizing binary cross-entropy and InfoNCE losses to achieve effective vision-language alignment.
  • 4.
    fVLM [2]: This method utilizes fine-grained contrastive learning for anatomy-level alignment between CT images and radiology reports. It addresses false negatives with a dual reduction strategy and employs co-teaching to improve semantic representation and diagnostic accuracy.
@@
To comprehensively evaluate the zero-shot diagnostic performance of our proposed method across various abnormalities, we introduce two new validation datasets in addition to existing public benchmarks [15]: CT-RATE-LT, consisting of 15 types of abnormalities, and RAD-ChestCT-LT, which encompasses 56 previously untested categories from the RAD-ChestCT dataset. As shown in Table 2, we perform a comparative analysis of the performance of the open-source models and our proposed method on these two datasets. The results reveal that our method consistently outperforms other approaches in zero-shot diagnostic performance across a wide range of abnormalities. These improvements can be attributed to the effective vision-language alignment achieved by combining semantic summarization and CMKI module.
Table 3: Performance comparison for volume-to-volume retrieval task on CT-RATE and RAD-ChestCT datasets.
@@
CT-RATE CT-Net 59.4 48.1 40.7
VocabFine 68.3 57.2 48.8
ClassFine 67.9 56.8 48.5
CT-CLIP 68.3 57.2 48.9
Merlin 62.6 51.3 43.9
BrgSA 69.2 58.5 50.5
RAD-ChestCT CT-Net 71.2 58.9 52.4
VocabFine 73.8 63.1 56.4
ClassFine 73.8 63.1 56.3
CT-CLIP 72.3 60.9 53.8
Merlin 74.0 63.3 56.6
BrgSA 74.3 63.8 57.2
Table 4: Performance comparison for report-to-volume retrieval task on CT-RATE dataset.
@@
VocabFine 0.1 0.6 2.3 2.0
CT-CLIP 2.9 5.0 18.0 28.7
Merlin 1.5 2.7 7.7 12.7
BrgSA 5.8 10.1 28.6 42.0
@@ 4.6 Volume-to-Volume Retrieval Task
As shown in Table 3, our method achieves superior performance, which can be attributed to the cross-modal interaction module proposed in this study. By constructing the CMKB as a bridge for image-text features, our method effectively projects features into a shared space, significantly reducing the discrepancies between the image and text modalities. Furthermore, contrastive learning is employed to align the image and text features, allowing the model to capture highly discriminative semantic information from images. As a result, our method delivers outstanding performance in the volume-to-volume retrieval task.
4.7 Report-to-Volume Retrieval Task
The report-to-volume retrieval task matches textual descriptions in radiology reports with their corresponding 3D medical imaging volumes. Retrieval is performed by calculating cosine similarity between the text and image embeddings, ranking volumes based on similarity scores, and selecting the top-$P$ most relevant ones. Since the RAD-ChestCT dataset does not include radiology reports, this evaluation is conducted on the CT-RATE dataset.
As shown in Table 4, VocabFine uses abnormality labels for fine-tuning. The training process retains some degree of vision-language alignment, but the images are only aligned with the labels and fail to achieve semantic alignment with the complete reports. This limitation results in the poorest performance in report-to-volume retrieval tasks. On the other hand, Merlin employs structured annotations to pre-train the vision encoder before aligning images and text. However, this strategy introduces bias into the image features, making effective semantic alignment difficult, which leads to poor retrieval results between images and reports. In contrast, our method leverages a large language model to semantically summarize the reports, combining the original reports and their semantic summaries during training. This approach reduces the textual training burden while preserving the rich semantic information in the reports. Our proposed BrgSA constructs a bridge between image and text information, enabling effective semantic alignment. Compared to CT-CLIP, our method achieves outstanding performance in the report-to-volume retrieval task, as evidenced by significant improvements in Recall@10 (5.0 vs. 10.1) and Recall@100 (28.7 vs. 42.0).
Table 5: Ablation study on the Semantic Summarization (SS) and the CMKI module for zero-shot abnormality diagnosis on the RAD-ChestCT dataset.
SS CMKI AUC ACC F1 score Precision
65.0 61.1 65.4 35.7
67.9 64.0 67.9 37.2
68.0 64.0 67.9 38.0
70.0 65.5 69.3 39.1
Table 6: Ablation study on the impact of the Semantic Summarization (SS) and the CMKI module for the report-to-volume retrieval task on the CT-RATE validation dataset.
SS CMKI Recall@5 Recall@10 Recall@50 Recall@100
2.2 3.7 13.1 22.2
2.2 4.1 18.4 29.1
3.4 5.8 21.8 34.6
5.8 10.1 28.6 42.0
5 Discussion
Fig. 6: Results of varying $K$ for CMKB on RAD-ChestCT dataset.
Fig. 7: UMAP visualization of features. Cosine similarity is used to evaluate the alignment between image and text features. The text features are generated using generic descriptive texts to ensure that all images can be matched to all texts. (a)–(c) represent features trained by BrgSA.

### v2

@@ 4.1 Materials
We conduct experiments on three publicly available datasets: CT-RATE [12], RAD-ChestCT [18], and INSPECT [42]. Detailed descriptions of these datasets are provided below.
CT-RATE [12]: The CT-RATE dataset comprises 25,692 non-contrast chest CT volumes, expanded to 50,188 volumes through various reconstruction techniques, originating from 21,304 unique patients. Each CT volume is paired with a corresponding radiology report. A total of 47,149 CT volumes with corresponding reports are used for training, while 3,039 CT volumes are reserved for internal validation. The dataset includes 18 abnormality labels, extracted from reports using an automated algorithm.
To further evaluate the zero-shot detection capabilities of our method, we construct an extended CT-RATE-LT dataset focusing on long-tail abnormalities. Specifically, we first predefine a set of 15 underrepresented abnormalities based on their clinical relevance and low prevalence in CT-RATE. GPT-4 Turbo  [^1] https://openai.com/index/gpt-4/ is then prompted to determine the presence or absence of these abnormalities in the impression sections of the reports. All GPT-4 annotations are subsequently verified by a board-certified radiologist, who carefully compares each predicted label with the original report description. Inconsistent or ambiguous cases are corrected based on clinical context, ensuring that each abnormality label precisely reflects the radiology report. The frequency distribution of CT-RATE-LT abnormalities is shown in Fig. 4.
RAD-ChestCT [18]: The RAD-ChestCT dataset contains 3,630 CT volumes, uniformly reconstructed using a single technique. This dataset includes 84 abnormality labels. As part of our external validation, as shown in Fig. 5, 27 abnormalities from RAD-ChestCT are mapped to the 18 abnormalities in CT-RATE to ensure consistency. To further demonstrate the strength of our method, we refer to the remaining 56 abnormalities from RAD-ChestCT as RAD-ChestCT-LT [18] and evaluate the zero-shot diagnosis performance on these additional abnormalities. Note that the abnormality, “other_path”, is excluded from our analysis due to its lack of meaningful semantic descriptions.
INSPECT [42]: We use the official test set comprising 3,214 contrast-enhanced CT pulmonary angiography (CTPA) cases to evaluate model generalization on the pulmonary embolism (PE) diagnosis task.
@@ 4.2 Dataset Preprocessing
For the image data, we unify CT volume spacing to $1.5\times 1.5\times 3\,\mathrm{mm}$ for uniform resolution across datasets. Random cropping during training generates input volumes of shape $224\times 224\times 112$, balancing computational efficiency and spatial coverage for 3D images. The ViT-Base patch size is $16\times 16\times 8$ for 3D inputs. To enhance feature representation, CT images are clipped to [1000,1000][-1000,1000] HU to focus on clinically relevant tissues and normalized to [1,1][-1,1].
For the text data, we process both the original reports and the semantic summaries generated by GPT-4 Turbo via the OpenAI API without additional fine-tuning, splitting them into shorter sentences for analysis. As part of our data augmentation strategy, five sentences are randomly sampled and concatenated into a longer sentence during training, enabling diverse text representations that better capture the variability in medical reporting styles.
Table 1: Zero-shot abnormality diagnosis performance comparison across internal (CT-RATE) and external (RAD-ChestCT) validation benchmarks. indicates that statistical significance ($p<0.001$) is confirmed for BrgSA against other methods.
Method CT-RATE RAD-ChestCT
AUC ACC F1 Pre mAP Rec@1 Pre@3 AUC ACC F1 Pre mAP Rec@1 Pre@3
[…]
BIUD [13] 71.3 68.1 71.6 33.8 62.9 60.6 65.2 33.7
[…]
X2CT-CLIP [44] 71.6 64.5
[…]
@@ 4.3 Implementation Details
In our experiments, we adapt ViT-B/16 [39] as the image encoder by replacing the 2D convolutions in the patch-embedding module with 3D convolutions, yielding a 3D ViT-B/16. For the text encoder, we fine-tune CXRBERT [40] on CT-RATE reports to capture medical-domain language. We also use M3AE [45] to pretrain both encoders. M3AE employs masked cross-modal reconstruction, where masked regions in one modality are reconstructed from the other, enhancing image–text semantic consistency and alignment. The CMKB basis vectors are initialized using orthogonal initialization.
In our experiments, $\alpha$, $\beta$, and $\gamma$ are empirically set to 0.5, 1, and 1, respectively, based on preliminary trials to balance the contributions of different loss components. $K$ is set to 2048, as determined through the experimental search results presented later. We use the Adam optimizer with a learning rate of $5\times 10^{-5}$. For the ablation studies, the batch size is set to 8 due to computational cost considerations, while our final configuration adopts a batch size of 64, which yields the best overall performance. The code is implemented using the PyTorch framework, and all experiments are conducted on a single NVIDIA A800 GPU. During the inference phase, we perform zero-shot abnormality diagnosis by setting the prompt template as “There is [abnormality]” and computing the similarity score for each category as
$s_{c}=\tfrac{1}{2}\big(\text{sim}(V,T_{c})+\text{sim}(\hat{V},T_{c})\big).$
where $s_{c}$ denotes the final similarity score between the image and category $c$. For cross-modal retrieval tasks, only the original reports are used as queries, and no additional summarization step is required, ensuring efficient inference.
@@ 4.4 Evaluation Metrics
For the multi-label classification, we evaluate zero-shot abnormality diagnosis using Area under the ROC Curve (AUC), Accuracy (ACC), F1 score (F1), and Precision (Pre). To better capture clinical usefulness, we additionally report mean Average Precision (mAP), Recall@1 (Rec@1), and Precision@3 (Pre@3), where Recall@1 and Precision@3 reflect report-level retrieval performance. Following fVLM [15], we determine the threshold for F1 and Precision by searching the value that minimizes the Euclidean distance to the ideal point (FPR=0, TPR=1) on the ROC curve, representing the optimal trade-off between sensitivity and specificity. For the volume-to-volume retrieval task, we use Mean Average Precision at $Q$ (MAP@$Q$) metrics, where $Q$ represents the number of top-ranked retrieved volumes considered. In our experiment, $Q$ is set to $\{5,10,50\}$. For the report-to-volume retrieval task, we use Recall@$P$, where $P$ represents the number of retrieved images, to measure the proportion of correctly matched images within the top-$P$ retrieved results for a given query text. In our experiment, $P$ is set to $\{5,10,50,100\}$. To assess statistical significance, we apply per-class DeLong tests for AUC, paired bootstrap resampling for ACC, F1, Precision, and mAP, and paired $t$-tests for Rec@1 and Pre@3. For retrieval metrics (MAP@$Q$ and Recall@$P$), statistical significance is assessed over queries using paired bootstrap. Throughout the tables, indicates that statistical significance ($p<0.001$) is confirmed for BrgSA against other methods.
Table 2: Performance comparison for zero-shot abnormality diagnosis on CT-RATE-LT and RAD-ChestCT-LT datasets.
Method CT-RATE-LT RAD-ChestCT-LT
AUC ACC F1 Pre mAP Rec@1 Pre@3 AUC ACC F1 Pre mAP Rec@1 Pre@3
[…]
Table 3: Performance comparison for zero-shot diagnosis on the INSPECT dataset.
Method AUC ACC F1 Pre mAP
[…]
5 Results and Analysis
@@ 5.1 Comparison with State-of-the-art Methods
We compare our method against six zero-shot approaches, supervised learning, and fine-tuning methods, as summarized in Table 1. The evaluation is conducted on both internal (CT-RATE [12]) and external (RAD-ChestCT [18]) benchmarks for zero-shot abnormality diagnosis. To provide an evaluation, we detail the comparison methods as follows. Some methods do not report certain metrics in their original papers; we denote these missing values with “–” in the table.
Supervised Learning. CT-Net [18] uses a deep convolutional neural network with transfer learning and 3D convolutions to perform multi-disease classification.
Fine-tuning Methods. VocabFine [12] fine-tunes only the projection layers of the image and text encoders in CT-CLIP while freezing all other pre-trained layers, preserving open-vocabulary capabilities. ClassFine [12] adds and trains a new classification layer on CT-CLIP while optionally freezing the pre-trained layers to retain existing feature representations.
Zero-shot Methods. CT-CLIP [12] employs a contrastive vision-language pretraining approach to align 3D chest CT volumes with radiology reports, utilizing a vision transformer for 3D CT and a text transformer for embedding generation. BIUD [13] aligns 3D CT images with 2D X-ray images through a language-guided retrieval strategy and enhances alignment using contrastive learning and entity-focused masking. Merlin [14] employs multi-task learning to process 3D abdominal CT volumes by integrating structured electronic health record data and unstructured radiology reports, optimizing binary cross-entropy and InfoNCE losses to achieve effective vision-language alignment. DCFormer [43] proposes a lightweight 3D vision-language encoder using decomposed convolutions for efficient volumetric representation learning; integrated into a CLIP framework, it reduces computational cost while maintaining strong alignment between CT images and radiology reports. X2CT-CLIP [44] introduces a tri-modal contrastive learning framework that aligns CXRs with 3D CT volumes and CT reports in a shared latent space, transferring knowledge from pretrained CT encoders to a CXR encoder using simulated image triplets to enable CT-level multi-abnormality detection from CXRs. fVLM [15] utilizes fine-grained contrastive learning for anatomy-level alignment between CT images and radiology reports, addressing false negatives with a dual reduction strategy and employing co-teaching to improve semantic representation.
@@
To comprehensively evaluate the zero-shot diagnostic performance of our proposed method across various abnormalities, we introduce two new validation datasets in addition to existing public benchmarks: CT-RATE-LT, consisting of 15 types of abnormalities, and RAD-ChestCT-LT, which encompasses 56 previously untested categories from the RAD-ChestCT dataset and naturally constitutes an open-set scenario. As shown in Table 2, we perform a comparative analysis of the performance of the open-source models and our proposed method on these two datasets. The results reveal that our method consistently outperforms other approaches in zero-shot diagnostic performance across a wide range of abnormalities. This demonstrates that our framework generalizes well to unseen abnormalities under open-set conditions. We further validate on the INSPECT test set with contrast-enhanced CTPA cases for PE diagnosis in Table 3, providing an out-of-distribution evaluation. These improvements can be attributed to the effective vision-language alignment achieved by combining semantic summarization and CMKI module.
Table 4: Volume-to-volume retrieval performance.
@@
[…]
Table 5: Cross-modal retrieval performance on CT-RATE.
@@
Report-to-Volume Retrieval
[…]
X2CT-CLIP [44] 4.8 7.7
[…]
Volume-to-Report Retrieval
[…]
@@ 5.2 Volume-to-Volume Retrieval Task
As shown in Table 4, our method achieves superior performance, which can be attributed to the cross-modal interaction module proposed in this study. By constructing the CMKB as a bridge for image-text features, our method effectively projects features into a shared space, significantly reducing the discrepancies between the image and text modalities. As a result, our method delivers outstanding performance in the volume-to-volume retrieval task.
5.3 Cross-modal Retrieval Tasks
Cross-modal retrieval is conducted in two directions: report-to-volume and volume-to-report. Unlike the training phase where semantic summarization is introduced to facilitate alignment, here retrieval is conducted directly using the original reports. Retrieval is based on cosine similarity between text and image embeddings, where candidates are ranked and the top-$P$ volumes or reports are selected. Since RAD-ChestCT does not include radiology reports, all evaluations are carried out on the CT-RATE dataset.
As shown in Table 5, our method achieves the best performance across all metrics, substantially surpassing CT-CLIP, the strongest baseline. For report-to-volume retrieval, Recall@10 improves from 5.2 to 22.2 and Recall@100 from 28.3 to 56.9; for volume-to-report retrieval, Recall@10 rises from 4.7 to 25.5 and Recall@100 from 28.4 to 61.0. These gains stem from the proposed CMKB, which bridges image and text features in a shared space, reducing modality gaps and improving retrieval alignment.